OpenStudy (ledah):

What is the slope of a line perpendicular to x - 5y = -10?

OpenStudy (anonymous):

maybe if you have something like <1,-5>*<x,y>=0 once you find x and y just put it in form y/x maybe

OpenStudy (anonymous):

1*x+-5*y=0

OpenStudy (ledah):

its either -5 or -4

OpenStudy (anonymous):

i remember a negative inverse relationship so here is a better way solve for you to get it in y=mx+b form then take m and do this -1/m

OpenStudy (anonymous):

y = (1/5)*x+2 so it be -1/(1/5)=-5

OpenStudy (ledah):

oh. talk about a walkthrough.

OpenStudy (ledah):

thanks

OpenStudy (anonymous):

I need a walkthru for zelda :P

OpenStudy (ledah):

which one?

OpenStudy (anonymous):

all of them :P

OpenStudy (anonymous):

specifically the one for zelda a link to the pass in the snes but I am j/k I know where to find them. :P

OpenStudy (ledah):

I beat all of them o-o

OpenStudy (anonymous):

I only beaten on in the gbc zelda dx or something lol

OpenStudy (ledah):

ah.

OpenStudy (anonymous):

I could have beaten zelda on snes but I got stuck as soon as i was sent to the past lol.

OpenStudy (ledah):

lol happens.

OpenStudy (anonymous):

Ever played golden sun?